# Measurement-driven quantum advantages in shallow circuits

### v2

@@ Abstract
In this work, we suggest a different approach using mid-circuit measurements, which not only overcomes the above challenges but is also independently interesting. Basically, in our work, we show that constant-depth quantum circuits are substantially more powerful computationally when equipped with mid-circuit measurements compared to the unitary setting. This regime of noisy intermediate scale quantum (NISQ) devices with mid-circuit measurements has been dubbed NISQ+ Quek (2024): We present a NISQ+/NISQ computational separation under plausible assumptions, highlighting the computational power of measurements along the way. To put this into context, dynamic circuits with measurements and feed-forward have shown great potential in implementing certain deep quantum circuits and preparing highly entangled states Skinner et al. (2019); Piroli et al. (2021); Lu et al. (2022); Deshpande et al. (2024); DeCross et al. (2023); Foss-Feig et al. (2023); Buhrman et al. (2024); Iqbal et al. (2024); Zi et al. (2025); Bene Watts et al. (2025); McGinley et al. (2025); Yan et al. (2025); Bäumer et al. (2025, 2024); Baumer and Woerner (2024). By conditioning operations on measurement outcomes, these circuits create non-local “shortcuts” that enable the rapid generation of complex, long-range entanglement at minimal depth Lieb and Robinson (1972); Eisert and Osborne (2006); Bravyi et al. (2006); Bäumer et al. (2024); Baumer and Woerner (2024). Inspired by these advances, we introduce a measurement-driven “fan-out staircase” architecture that implements the dense, long-range interactions required for IQP computational hardness at constant depth, even on simple two-dimensional (2D) hardware. Importantly, we show that the output distributions from these shallow circuits retain key signatures of complexity (such as anti-concentration), providing strong evidence for their classical hardness while significantly relaxing the experimental requirements of depth and connectivity.
@@
where $\eta$ is a geometry-dependent constant. $\xi$ provides a lower bound for the geometrically local circuit cost required to prepare a given state from a product in two dimensions (see the SM). We benchmark against $\xi_{\text{lin}}$, the average for output states of deep, linear-depth random CX circuits ($\sim 6n$). Although such constant-depth circuits on 2D lattices are not expected to achieve genuine volume-law entropy across each cut—particularly for larger systems—Fig. 3(a) shows that measurement-driven IQP circuits nonetheless achieve $\xi/\xi_{\mathrm{lin}}\approx 1$, matching the entangling power of linear-depth CX circuits. In contrast, conventional depth-matched IQP circuits without auxiliary systems fall significantly short.
The collision probability, $\chi:={\mathbb{E}}_{\mathcal{C}}\left[\sum_{x}p_{\mathcal{C}}(x)^{2}\right]$, quantifies quantum circuit anti-concentration. For sufficiently deep Haar-random circuits, the collision probability approaches $\chi_{\text{Haar}}=2/(2^{n}+1)$. Anti-concentration requires the ratio $\chi/\chi_{\text{Haar}}$ to remain bounded by a small constant Dalzell et al. (2022), a condition underpinning complexity arguments and benchmarks like XEB Boixo et al. (2018); Arute and others (2019). Fig. 3(b) compares collision probabilities between measurement-driven IQP circuits and standard IQP circuits without auxiliary qubits at comparable depths. Unlike standard IQP circuits, measurement-driven IQP circuits exhibit a decreasing collision probability ratio with increasing system size, indicating robust anti-concentration at constant depth and minimal sensitivity to connectivity constraints.
@@
Figure 3: Comparison of measurement-driven IQP sampling (blue) with standard IQP sampling (red/gray) at matched constant two-qubit depth on a 2D nearest-neighbor grid. (a) Relative entanglement-based circuit cost $\xi/\xi_{\text{lin}}$ for fixed-depth circuits $\mathfrak{L}=2,\mathfrak{D}=2$ versus system size. (b) Collision probability ratio $\chi/\chi_{\text{Haar}}$ versus system size. (c) Total variation distance between noisy and ideal output distributions under gate depolarizing errors. (d) Total variation distance under pure dephasing noise assuming fixed gate duration.
@@
We compare four reservoirs: (i) a random anisotropic Heisenberg reservoir, $\mathcal{H}_{\mathrm{H}}=\sum_{\langle i,j\rangle}\sum_{\mu}J_{i,j}^{\mu}\,\sigma^{\mu}_{i}\sigma^{\mu}_{j}$, where $\mu\in\{x,y,z\}$ labels spin components; (ii) a random transverse-field Ising reservoir, $\mathcal{H}_{\mathrm{TFI}}=\sum_{\langle i,j\rangle}J_{i,j}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{x}_{i}$; (iii) a random transverse-field XY reservoir $\mathcal{H}_{\mathrm{XY}}=\sum_{\langle i,j\rangle}(J^{x}_{i,j}\sigma^{x}_{i}\sigma^{x}_{j}+J^{y}_{i,j}\sigma^{y}_{i}\sigma^{y}_{j})+\sum_{i}h_{i}\sigma^{x}_{i}$; and (iv) a measurement-based multibody XY reservoir $\mathcal{H}_{\mathbf{A}}=\sum_{i=1}^{s}(c^{x}_{i}\prod_{j\in A_{i}}\sigma^{x}_{j}+c^{y}_{i}\prod_{j\in A_{i}}\sigma^{y}_{j})$, with all coupling coefficients $J_{i,j}^{x,y,z}$, $h_{i}$, and $c_{i}^{x,y,z}$ drawn i.i.d. from $\mathcal{N}(0,1/2)$, $s=16$, and $\mathbf{A}$ encodes the measurement‑enabled architecture. All reservoirs use 16 system qubits on 2D-square and hexagonal lattices, undergo ten identical Floquet cycles generated by their Hamiltonian terms. The classification of phases is performed using an SVM with an RBF kernel Cortes and Vapnik (1995); Chang and Lin (2011).
@@
$\bigl\lvert\operatorname{Tr}\bigl[O\,\mathcal{R}_{\mathrm{MD}}^{(T)}(\lvert\psi_{0}\rangle)\bigr]-\operatorname{Tr}\bigl[O\,\mathcal{R}_{\mathrm{MD}}^{(T)}(\lvert\psi_{1}\rangle)\bigr]\bigr\rvert=\Omega(1),$ (3)
@@ Acknowledgements.
We thank Antonio Anna Mele, Marcel Hinsche, Elies Gil-Fuster, Zhenhuan Liu, Julio Carlos Magdalena De La Fuente, and Xiaoyu He for insightful discussions and valuable feedback. This work has been supported by the BMFTR (DAQC, MuniQC-Atoms), Clusters of Excellence (ML4Q, MATH+), the Munich Quantum Valley, Berlin Quantum, the Quantum Flagship (MILLENION, PASQUANS2), the DFG (CRC 183, SPP 2541), the European Research Council (DebuQC), and the Alexander-von-Humboldt Foundation.
@@ A.0.1 One-dimensional topology
with each auxiliary qubit initially prepared in the state vector $|0\rangle$. We denote the controlled-NOT (CX) gate with control qubit $Q_{i}$ and target qubit $\widetilde{Q}_{j}$ as $\operatorname{CX}(Q_{i},\widetilde{Q}_{j})$, and the reverse operation, where $\widetilde{Q}_{j}$ is the control and $Q_{i}$ is the target, as $\operatorname{CX}(\widetilde{Q}_{j},Q_{i})$.
@@
$U^{(1)}=\left(\prod_{i=1}^{n-1}\operatorname{CX}(\widetilde{Q}_{i},Q_{i+1})\right)\cdot\left(\prod_{i=1}^{n-1}\operatorname{CX}(Q_{i},\widetilde{Q}_{i})\right).$ (S3)
We now show that if, for a given auxiliary qubit $\widetilde{Q}_{i}$ (initialized in $|0\rangle$), we sequentially apply $\operatorname{CX}(Q_{i},\widetilde{Q}_{i})$ and $\operatorname{CX}(\widetilde{Q}_{i},Q_{i+1})$, then subsequently applying a Hadamard gate on $\widetilde{Q}_{i}$ and measuring $\widetilde{Q}_{i}$ in the computational basis with outcome $|0\rangle$ yields an effective operation on the system qubits equivalent to a $\mathrm{CX}$ gate from $Q_{i}$ to $Q_{i+1}$. For clarity, we now detail the state evolution.
@@
The application of the first gate, $\operatorname{CX}(Q_{i},\widetilde{Q}_{i})$, transforms the state vector into
@@
Subsequently, the application of the second gate, $\operatorname{CX}(\widetilde{Q}_{i},Q_{i+1})$, results in the state vector
@@
$Z(\widetilde{Q}_{j})\left(\prod_{i}^{n-1}\operatorname{CX}(Q_{i},\widetilde{Q}_{i})\right)$ (S8)
$=$ $\left(\prod_{i}^{n-1}\operatorname{CX}(Q_{i},\widetilde{Q}_{i})\right)Z(Q_{j})Z(\widetilde{Q}_{j}),$
@@
$Z(\widetilde{Q}_{j-1})Z(Q_{j})\left(\prod_{i=1}^{n-1}\operatorname{CX}(\widetilde{Q}_{i},Q_{i+1})\right)$ (S9)
$=$ $\left(\prod_{i=1}^{n-1}\operatorname{CX}(\widetilde{Q}_{i},Q_{i+1})\right)Z(Q_{j}),$
@@
$\operatorname{CX}(Q_{i},\widetilde{Q}_{i}),\quad\operatorname{CX}(Q_{i+1},\widetilde{Q}_{i+1}),$ (S14)
$\operatorname{CX}(\widetilde{Q}_{i},Q_{i+1}),\quad\operatorname{CX}(\widetilde{Q}_{i+1},Q_{i+2}),$
$\operatorname{CX}(Q_{i+1},\widetilde{Q}_{i+1}),\quad\operatorname{CX}(\widetilde{Q}_{i+1},Q_{i+2}).$
@@
Thus, obtaining a measurement outcome $m_{j}=1$ on $\widetilde{Q}_{j}$ is equivalent to the ideal outcome $m_{j}=0$ up to the application of corrective $Z$ operations on the system qubits, specifically on those qubits whose indices belong to $I_{2}(j)$. The transfer matrix here have entries
@@ Appendix B Appendix B: Measurement-driven IQP sampling
$Z=(I-2|1\rangle\langle 1|)=I-2|1\rangle\langle 1|,$ (S35)
@@
We can express this expansion via binary strings $q\in\{0,1\}^{r}$. Specifically, let $(\frac{Z-I}{2})_{q}$ denote a tensor product of $r$ factors where each factor is $(\frac{Z-I}{2})$ is $q_{j}=1$ and I otherwise (for position $j\in\{1,\dots,r\}$. Equivalently,
@@
Rewriting each $(\frac{Z-I}{2})_{q}$ as a tensor product of $I$ and $Z$, we see that each surviving term of Hamming weight $|q|\leq k$ is effectively a $|q|$-body $Z$-interaction. In total, the exponent can be broken down into a linear combination of diagonal operators of body-size at most $k$. Hence, $U_{r}(\ell,k)$ can be decomposed into a product of multi-qubit diagonal operators, each of which involves at most $k$ qubits, qubits, up to an irrelevant global phase.
@@
The measurement-driven IQP circuits $\mathcal{C}_{\mathsf{FS}}$ interleaves fan-out staircases with parallel single-qubit diagonal gates $\exp(\mathrm{i}\vartheta_{i,j}Z_{j})$ for preparing Hamiltonian phase states. Any effective multi-qubit diagonal rotation in $\mathcal{C}_{\mathsf{FS}}$ ultimately reduces to a product of multi-qubit diagonal unitaries of the form
@@
From Lemma 1, each such $r$-qubit rotation with angle $\ell\pi/2^{k}$ is equivalent (up to a global phase) to combining at most $k$-body diagonal $Z$-interactions with integer multiples of $2\pi$. Hence, when all angles $\bm{\vartheta}$ lie in $\{\ell\,\pi/2^{k}\}$, $\mathcal{C}_{\mathsf{FS}}$ can be fully synthesized by commuting $Z$-interactions of at most $k$ qubits each. Because all gates remain diagonal in the computational basis, they commute pairwise. The net effect is equivalent to a $k$-local IQP with commuting $Z$-gates of weight at most $k$.

### v3

@@ Abstract
A closely related precedent is the quench-based architecture of Bermejo-Vega et al. Bermejo-Vega et al. (2018), which realizes constant-depth hard sampling on two-dimensional nearest-neighbor lattices using product-state inputs, short-time evolution under a translation-invariant Ising Hamiltonian, and terminal fixed-basis readout—all without mid-circuit measurements. That scheme yields families of nearest-neighbor, translation-invariant 2-local constant-depth IQP circuits, with the complexity-theoretic route proceeding through a nonadaptive measurement-based quantum computing encoding of deeper logical circuits on an enlarged two-dimensional lattice. Our goal here is complementary: we ask whether mid-circuit measurements and feed-forward can compress a class of dense, long-range IQP-type diagonal circuits to constant depth on bounded-degree hardware, using auxiliary qubits that are measured out and do not appear in the final output register.
In this work, we suggest a different approach using mid-circuit measurements, which not only addresses the depth and connectivity challenges facing standard IQP implementations but is also independently interesting. Basically, in our work, we show that constant-depth quantum circuits are substantially more powerful computationally when equipped with mid-circuit measurements compared to the unitary setting. This regime of noisy intermediate scale quantum (NISQ) devices with mid-circuit measurements has been dubbed NISQ+ Quek (2024): We present a NISQ+/NISQ computational separation under plausible assumptions, highlighting the computational power of measurements along the way. To put this into context, dynamic circuits with measurements and feed-forward have shown great potential in implementing certain deep quantum circuits and preparing highly entangled states Skinner et al. (2019); Piroli et al. (2021); Lu et al. (2022); Deshpande et al. (2024); DeCross et al. (2023); Foss-Feig et al. (2023); Buhrman et al. (2024); Iqbal et al. (2024); Zi et al. (2025); Bene Watts et al. (2025); McGinley et al. (2025); Yan et al. (2025); Bäumer et al. (2025, 2024); Baumer and Woerner (2024). By conditioning operations on measurement outcomes, these circuits create non-local “shortcuts” that enable the rapid generation of complex, long-range entanglement at minimal depth Lieb and Robinson (1972); Eisert and Osborne (2006); Bravyi et al. (2006); Bäumer et al. (2024); Baumer and Woerner (2024). Inspired by these advances, we introduce a measurement-driven “fan-out staircase” architecture that implements the dense, long-range interactions required for IQP computational hardness at constant depth, even on simple two-dimensional (2D) hardware. Importantly, we show that the output distributions from these shallow circuits retain key signatures of complexity (such as anti-concentration), providing strong evidence for their classical hardness while significantly relaxing the experimental requirements of depth and connectivity.
@@
where $\eta$ is a geometry-dependent constant. $\xi$ provides a lower bound for the geometrically local circuit cost required to prepare a given state from a product in two dimensions (see the SM). We benchmark against $\xi_{\mathrm{lin}}$, the average for output states of deep, linear-depth random CX circuits ($\sim 6n$). Although such constant-depth circuits on 2D lattices are not expected to achieve genuine volume-law entropy across each cut—particularly for larger systems—Fig. 3(a) shows that measurement-driven IQP circuits nonetheless achieve $\xi/\xi_{\mathrm{lin}}\approx 1$, matching the entangling power of linear-depth CX circuits. In contrast, conventional depth-matched IQP circuits without auxiliary systems fall significantly short.
The collision probability, $\chi:={\mathbb{E}}_{\mathcal{C}}\left[\sum_{x}p_{\mathcal{C}}(x)^{2}\right]$, quantifies quantum circuit anti-concentration. For sufficiently deep Haar-random circuits, the collision probability approaches $\chi_{\text{Haar}}=2/(2^{n}+1)$. Anti-concentration requires the ratio $\chi/\chi_{\text{Haar}}$ to remain bounded by a small constant Dalzell et al. (2022), a condition underpinning complexity arguments and benchmarks like cross-entropy benchmarking Boixo et al. (2018); Arute and others (2019). Fig. 3(b) compares collision probabilities between measurement-driven IQP circuits and standard IQP circuits without auxiliary qubits at comparable depths. Unlike standard IQP circuits, measurement-driven IQP circuits exhibit a decreasing collision probability ratio with increasing system size, indicating robust anti-concentration at constant depth and minimal sensitivity to connectivity constraints.
@@
Figure 3: Comparison of measurement-driven IQP sampling (blue) with standard IQP sampling (red/gray) at matched constant two-qubit depth on a 2D nearest-neighbor grid. (a) Relative entanglement-based circuit cost $\xi/\xi_{\mathrm{lin}}$ for fixed-depth circuits $\mathfrak{L}=2,\mathfrak{D}=2$ versus system size. (b) Collision probability ratio $\chi/\chi_{\text{Haar}}$ versus system size. (c) Total variation distance between noisy and ideal output distributions under gate depolarizing errors. (d) Total variation distance under pure dephasing noise assuming fixed gate duration.
@@
We compare four reservoirs: (i) a random anisotropic Heisenberg reservoir, $\mathcal{H}_{\mathrm{H}}=\sum_{\langle i,j\rangle}\sum_{\mu}J_{i,j}^{\mu}\,\sigma^{\mu}_{i}\sigma^{\mu}_{j}$, where $\mu\in\{x,y,z\}$ labels spin components; (ii) a random transverse-field Ising reservoir, $\mathcal{H}_{\mathrm{TFI}}=\sum_{\langle i,j\rangle}J_{i,j}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i}h_{i}\sigma^{x}_{i}$; (iii) a random transverse-field XY reservoir $\mathcal{H}_{\mathrm{XY}}=\sum_{\langle i,j\rangle}(J^{x}_{i,j}\sigma^{x}_{i}\sigma^{x}_{j}+J^{y}_{i,j}\sigma^{y}_{i}\sigma^{y}_{j})+\sum_{i}h_{i}\sigma^{x}_{i}$; and (iv) a measurement-based multibody XY reservoir $\mathcal{H}_{\mathbf{A}}=\sum_{i=1}^{s}(c^{x}_{i}\prod_{j\in A_{i}}\sigma^{x}_{j}+c^{y}_{i}\prod_{j\in A_{i}}\sigma^{y}_{j})$, with all coupling coefficients $J_{i,j}^{x,y,z}$, $h_{i}$, and $c_{i}^{x,y,z}$ drawn i.i.d. from $\mathcal{N}(0,1/2)$, $s=16$, and $\mathbf{A}$ encodes the measurement‑enabled architecture. All reservoirs use 16 system qubits on 2D-square and hexagonal lattices, undergo ten identical Floquet cycles generated by their Hamiltonian terms. The classification of phases is performed using a support vector machine (SVM) with a radial basis function kernel Cortes and Vapnik (1995); Chang and Lin (2011).
@@
[…]
@@ Acknowledgements.
We thank Antonio Anna Mele, Marcel Hinsche, Elies Gil-Fuster, Zhenhuan Liu, Julio Carlos Magdalena De La Fuente, and Xiaoyu He for insightful discussions and valuable feedback. This work has been supported by the BMFTR (DAQC, MuniQC-Atoms), Clusters of Excellence (ML4Q, MATH+), the Munich Quantum Valley, Berlin Quantum, the Quantum Flagship (MILLENION, PASQUANS2), the DFG (CRC 183, SPP 2514), the European Research Council (DebuQC), and the Alexander-von-Humboldt Foundation.
@@ A.0.1 One-dimensional topology
with each auxiliary qubit initially prepared in the state vector $|0\rangle$. We denote the controlled-NOT (CX) gate with control qubit $Q_{i}$ and target qubit $\widetilde{Q}_{j}$ as $\mathrm{CX}(Q_{i},\widetilde{Q}_{j})$, and the reverse operation, where $\widetilde{Q}_{j}$ is the control and $Q_{i}$ is the target, as $\mathrm{CX}(\widetilde{Q}_{j},Q_{i})$.
@@
$U^{(1)}=\left(\prod_{i=1}^{n-1}\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})\right)\cdot\left(\prod_{i=1}^{n-1}\mathrm{CX}(Q_{i},\widetilde{Q}_{i})\right).$ (S3)
We now show that if, for a given auxiliary qubit $\widetilde{Q}_{i}$ (initialized in $|0\rangle$), we sequentially apply $\mathrm{CX}(Q_{i},\widetilde{Q}_{i})$ and $\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})$, then subsequently applying a Hadamard gate on $\widetilde{Q}_{i}$ and measuring $\widetilde{Q}_{i}$ in the computational basis with outcome $|0\rangle$ yields an effective operation on the system qubits equivalent to a $\mathrm{CX}$ gate from $Q_{i}$ to $Q_{i+1}$. For clarity, we now detail the state evolution.
@@
The application of the first gate, $\mathrm{CX}(Q_{i},\widetilde{Q}_{i})$, transforms the state vector into
@@
Subsequently, the application of the second gate, $\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})$, results in the state vector
@@
$Z(\widetilde{Q}_{j})\left(\prod_{i}^{n-1}\mathrm{CX}(Q_{i},\widetilde{Q}_{i})\right)$ (S8)
$=$ $\left(\prod_{i}^{n-1}\mathrm{CX}(Q_{i},\widetilde{Q}_{i})\right)Z(Q_{j})Z(\widetilde{Q}_{j}),$
@@
$Z(\widetilde{Q}_{j-1})Z(Q_{j})\left(\prod_{i=1}^{n-1}\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})\right)$ (S9)
$=$ $\left(\prod_{i=1}^{n-1}\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1})\right)Z(Q_{j}),$
@@
$\mathrm{CX}(Q_{i},\widetilde{Q}_{i}),\quad\mathrm{CX}(Q_{i+1},\widetilde{Q}_{i+1}),$ (S14)
$\mathrm{CX}(\widetilde{Q}_{i},Q_{i+1}),\quad\mathrm{CX}(\widetilde{Q}_{i+1},Q_{i+2}),$
$\mathrm{CX}(Q_{i+1},\widetilde{Q}_{i+1}),\quad\mathrm{CX}(\widetilde{Q}_{i+1},Q_{i+2}).$
@@
Thus, obtaining a measurement outcome $m_{j}=1$ on $\widetilde{Q}_{j}$ is equivalent to the ideal outcome $m_{j}=0$ up to the application of corrective $Z$ operations on the system qubits, specifically on those qubits whose indices belong to $I_{2}(j)$. The transfer matrix here has entries
@@ Appendix B Appendix B: Measurement-driven IQP sampling
$Z=I-2|1\rangle\langle 1|,$ (S35)
@@
We can express this expansion via binary strings $q\in\{0,1\}^{r}$. Specifically, let $(\frac{Z-I}{2})_{q}$ denote a tensor product of $r$ factors where each factor is $(\frac{Z-I}{2})$ if $q_{j}=1$ and $I$ otherwise (for position $j\in\{1,\dots,r\}$. Equivalently,
@@
Rewriting each $(\frac{Z-I}{2})_{q}$ as a tensor product of $I$ and $Z$, we see that each surviving term of Hamming weight $|q|\leq k$ is effectively a $|q|$-body $Z$-interaction. In total, the exponent can be broken down into a linear combination of diagonal operators of body-size at most $k$. Hence, $U_{r}(\ell,k)$ can be decomposed into a product of multi-qubit diagonal operators, each of which involves at most $k$ qubits, up to an irrelevant global phase.
@@
The measurement-driven IQP circuits $\mathcal{C}_{\mathrm{MD}}$ interleave fan-out staircases with parallel single-qubit diagonal gates $\exp(\mathrm{i}\vartheta_{i,j}Z_{j})$ for preparing Hamiltonian phase states. Any effective multi-qubit diagonal rotation in $\mathcal{C}_{\mathrm{MD}}$ ultimately reduces to a product of multi-qubit diagonal unitaries of the form
@@
From Lemma 1, each such $r$-qubit rotation with angle $\ell\pi/2^{k}$ is equivalent (up to a global phase) to combining at most $k$-body diagonal $Z$-interactions with integer multiples of $2\pi$. Hence, when all angles $\bm{\vartheta}$ lie in $\{\ell\,\pi/2^{k}\}$, $\mathcal{C}_{\mathrm{MD}}$ can be fully synthesized by commuting $Z$-interactions of at most $k$ qubits each. Because all gates remain diagonal in the computational basis, they commute pairwise. The net effect is equivalent to a $k$-local IQP with commuting $Z$-gates of weight at most $k$.